OpenStudy (anonymous):

okay medal will be awarded to whoever gives the best answer. What polynomial has roots of –6, 1, and 4 A. x3 – 9x2 – 22x + 24 B. x3 – x2 – 26x – 24 C. x3 + x2 – 26x + 24 D. x3 + 9x2 + 14x – 24

Parth (parthkohli):

If \(\rm blah\) is a root, then \(x - \rm blah\) is a factor. We might know that a polynomial is just a product of factors.

Parth (parthkohli):

Can you write the three factors down? I'd write down one for ya. \(-6\) is a root, so \(x - (-6) = x + 6\) is a factor.

OpenStudy (anonymous):

yea but I cant figure it out

OpenStudy (anonymous):

x=6 x-1 x-4

Parth (parthkohli):

@Luis_Rivera Eh no...

Parth (parthkohli):

More like \((x + 6)(x -1)(x - 4)\)!

OpenStudy (anonymous):

hmm so now what

Parth (parthkohli):

Now just expand \((x + 6)\times (x - 1) \times (x - 4)\)

OpenStudy (anonymous):

so (x+6)(x-1) and that answer time (x-4)

OpenStudy (anonymous):

nope but I wouldnt mind some

OpenStudy (anonymous):

the expression which give the result zero when substituted with all the three roots

OpenStudy (anonymous):

so I multiply (x+6)(x-1) and got x2 + 5x - 6 now i multiply that by (x-4)?

OpenStudy (anonymous):

so the answers c

Parth (parthkohli):

@Luis_Rivera lol no!

Parth (parthkohli):

I mean, the factors are NOT \(x-6,x+1,x+4\) but their respective conjugates.